OpenStudy (anonymous):

If b(x) = 4x + 1, then what is the value of b(4)? a) 1 b) 9 c) 17 d) 20

OpenStudy (anonymous):

4(4) + 1 = 16+1 = 17 answer is C

OpenStudy (anonymous):

b(4)= 4*(4)+1